# Restyling Unsupervised Concept Based Interpretable Networks with Generative Models

### v1

@@ 0.D.0.3 Selecting number of concepts $K$
K  Accuracy  MSE  LPIPS  FID
@@
0.D.0.4 Orthogonality loss
For the final layer of $\Psi$, we choose a $1\times 1$ convolutional layer. Thus, the weights/kernels for this layer can be represented as single matrix of size number of input feature maps times number of concepts $K$. We encourage the $\ell_{2}$ normalized columns to be orthogonal which in turn encourages each $\phi_{k}(x)$ to be predicted using different feature maps. We report the quantitative effect of this loss in Tab. 9. Incorporating this loss offers slight advantage in improved perceptual similarity. However, another key reason we incorporated this loss in our experiments is that we qualitatively observed a greater propensity of multiple concepts highly relevant for a class to capture a common concept about that class. This loss thus offered a way to encourage different concepts to rely on different feature maps. Note that it only affects parameters of final layer of $\Psi$ that outputs $\Phi(x)$, and we do not use any additional hyperparameter for it.
@@ 0.D.0.4 Orthogonality loss
$\mathcal{L}_{orth}$  Accuracy  MSE  LPIPS  FID
@@
0.D.0.5 Other loss weigths
We report the accuracy and reconstruction metrics for different $\alpha$ (weight for output fidelity loss) and $\delta$ (weight for sparsity of activations). A small weight on output fidelity impacts the performance of the system and a high weight affects the reconstruction without benefiting the performance much. We found a balance with $\alpha=0.5$ which we employed for all datasets. A high $\delta$ impacts both the performance and reconstruction. For CUB-200 and Stanford Cars, due to prioritizing reconstruction we opted for a smaller $\delta=0.2$, while for CelebA-HQ, since obtaining a good reconstruction was relatively easier, we opted for a higher $\delta=2$.
@@ 0.D.0.5 Other loss weigths
$\alpha$  Accuracy  MSE  LPIPS  FID
@@
$\delta$  Accuracy  MSE  LPIPS  FID
@@
Appendix 0.E Additional visualization
@@
0.E.0.1 Use of difference images can be useful
@@
Appendix 0.F Limitations of approach
  • While the idea in principle should be adaptable to other choices of pretrained $G$, assuming other choices are inline with desiderata discussed in Sec. 0.A.1, it requires redesigning $\Omega$ accordingly. The current VisCoIN design in our experiments is only compatible with StyleGAN.
@@ Appendix 0.F Limitations of approach
  • While VisCoIN significantly upgrades the visualization pipeline over previous CoIN architectures, some of our choices, arising from restrictions to preserve a by-design interpretable predictive structure or limited computational resources, lead to certain limitations. (1) The choice of using a small dictionary as bottleneck under multiple other constraints leads to worse reconstruction compared to StyleGAN inversion architectures. (2) The choice of using a concept dictionary not defined in the Stylespace, results in learnt concepts that are not as well disentangled as Stylespace. Dictionaries for post-hoc interpretation defined in Stylespace provide more disentangled concepts for interpretation (as in StyleEx [43]), but are significantly larger and thus less suitable for by-design architecture. Finally, (3) the choice of using a pretrained $G$ improves training time, complexity and reusability, but also implies that the system’s quality is limited by the quality of the pretrained $G$. For instance, for visualization, if $G$ can’t generate some specific feature, it can be difficult to visualize a concept $\phi_{k}$ that encodes that feature.
  • For the case of single FC layers in $\Omega$ our visualization process follows linear trajectories in $\mathcal{W}^{+}$ latent space of StyleGAN when modifying an activation. Recent work has shown that linear trajectories are not necessarily optimal for latent traversals [65].
Appendix 0.G Potential negative impacts

### v2

@@ F.3 Selecting number of concepts $K$
K  Accuracy ($\uparrow$)  MSE ($\downarrow$)  LPIPS ($\downarrow$)  FID ($\downarrow$)
@@
F.4 Orthogonality loss
For the final layer of $\Psi$, we choose a $1\times 1$ convolutional layer. Thus, the weights/kernels for this layer can be represented as single matrix of size number of input feature maps times number of concepts $K$. We encourage the $\ell_{2}$ normalized columns to be orthogonal which in turn encourages each $\phi_{k}(x)$ to be predicted using different feature maps. We report the quantitative effect of this loss in Table 17. Incorporating this loss offers slight advantage in improved perceptual similarity. However, another key reason we incorporated this loss in our experiments is that we qualitatively observed a greater propensity of multiple concepts highly relevant for a class to capture a common concept about that class. This loss thus offered a way to encourage different concepts to rely on different feature maps. Note that it only affects parameters of final layer of $\Psi$ that outputs $\Phi(x)$, and we do not use any additional hyperparameter for it.
@@ F.4 Orthogonality loss
$\mathcal{L}_{orth}$  Accuracy ($\uparrow$)  MSE ($\downarrow$)  LPIPS ($\downarrow$)  FID ($\downarrow$)
@@
F.5 Other loss weigths
We report the accuracy and reconstruction metrics for different $\alpha$ (weight for output fidelity loss) and $\delta$ (weight for sparsity of activations). A small weight on output fidelity degrades the performance of the system and a high weight affects the reconstruction without benefiting the performance much. We found a balance with $\alpha=0.5$ which we employed for all datasets. A high $\delta$ impacts both the performance and reconstruction since it encourages activation of smaller number of concepts for any input more strongly. However, eliminating $\delta$ can result in poor sparsity and consequently interpretability of the concept activations for prediction. Our overall strategy was thus to use a high enough $\delta$ that it does not significantly affect reconstruction quality. For CUB-200 and Stanford Cars, due to a greater need of prioritizing reconstruction we opted for a smaller $\delta=0.2$, while for CelebA-HQ, since obtaining a good reconstruction was relatively easier, we opted for a higher $\delta=2$.
@@ F.5 Other loss weigths
$\alpha$  Accuracy ($\uparrow$)  MSE ($\downarrow$)  LPIPS ($\downarrow$)  FID ($\downarrow$)
@@
$\delta$  Accuracy ($\uparrow$)  MSE ($\downarrow$)  LPIPS ($\downarrow$)  FID ($\downarrow$)
@@
F.6 Baseline for concept translator $\Omega$
We compare our system with a variant where we directly use $\Phi(x)$ as latent vector $w_{x}$ for $G$, eliminating $\Omega$. While our model allows this design, it comes with certain limitations: (i) The user can’t control the number of concepts. They are forced to employ a concept dictionary of same size as dimension of the latent space. (ii) Since the generator is pretrained and fixed, the resulting $\Phi(x)$ learnt is not sparse. (iii) Finally, in particular for GANs, it forcibly associates concept functions with columns of identity matrix as directions in latent space. Using an $\Omega$ (for instance a linear layer) allows the model to learn general directions in the latent space to associate to each concept function, which aligns with the conventional strategy for latent traversal inside GAN. As can be seen in Table 20, removing $\Omega$ leads to poorer reconstruction.
[…]
$\Omega$  Accuracy ($\uparrow$)  MSE ($\downarrow$)  LPIPS ($\downarrow$)  FID ($\downarrow$)
Yes 79.44 0.161 0.545 15.85
No 79.24 0.182 0.572 15.96
F.7 Output fidelity loss
One can consider using cross-entropy loss with ground truth labels instead of “output fidelity loss”. We specify the possibility to use both when describing the general architecture of unsupervised CoINs (in Section 3.1). We decided to use the output fidelity loss following Sarkar et al. (2022). The current design also draws inspiration from knowledge distillation setting, in which a student model is trained to reproduce output of a teacher model. One additional perk of using this “output fidelity loss” during VisCoIN training is that it can also be applied for images without annotations, such as images sampled from $G$ (further details about VisCoIN training in Appendix C). This provides additional guidance and stability to train $g$. For completeness, we include in Table 21 an experiment of VisCoIN trained using a standard cross-entropy loss with ground truth labels on CUB-200 dataset.
[…]
Model Acc. $g$ ($\uparrow$) LPIPS ($\downarrow$) $FF_{x}$ ($\tau=0.2$) ($\uparrow$)
VisCoIN - Output fidelity loss 79.44 0.545 0.146
VisCoIN - Cross-entropy loss 78.89 0.559 0.076
Appendix G Additional Analysis
@@
Figure 11: Local interpretation on CUB dataset (sample from class 15) for different concepts along with their relevance scores.
@@
G.1 Local interpretations
We present in Fig. 11 local interpretations on test samples of CUB dataset, for different concepts relevant to that class. For a single test sample, a user obtains multiple concepts that are relevant for the prediction of the test sample. We extract the concepts with high relevances for the given input. For each concept one can use the visualization pipeline introduced to visualize what part of the input activates the respective concept. Note that the visualizations for $\phi_{110},\phi_{37}$ remain consistent with their global visualizations.
G.2 Use of difference images can be useful
@@
G.3 Average latent vector
We illustrate through example on CUB in Fig. 13 that the average latent vector of pretrained $G$ is typically representative of the dataset.
Figure 13: Generated image for average latent vector (“center” of latent space) in CUB (temporary figure).
Appendix H Limitations of approach
@@
  • The choice of using a pretrained $G$ improves training time, complexity and reusability, but also implies that the system’s quality is limited by the quality of the pretrained $G$. For instance, for visualization, if $G$ can’t generate some specific feature, it can be difficult to visualize a concept $\phi_{k}$ that encodes that feature.
  • For the case of single FC layers in $\Omega$, our visualization process follows linear trajectories in the latent space of $G$ when modifying an activation. Recent work has shown that linear trajectories are not necessarily optimal for latent traversals (Song et al., 2023).
Appendix I Potential negative impacts